# A Novel Large-scale Crop Dataset and Dual-stream Transformer Method for Fine-grained Hierarchical Crop Classification from Integrated Hyperspectral EnMAP Data and Multispectral Sentinel-2 Time Series

### v1

@@ Abstract
Fine-grained crop classification is crucial for precision agriculture and food security monitoring. It requires simultaneous capture of both phenological dynamics (obtained from multi-temporal satellite data like Sentinel-2) and subtle spectral variations (demanding nanometer-scale spectral resolution from hyperspectral imagery). Research combining these two modalities remains scarce currently due to challenges in hyperspectral data acquisition and crop types annotation costs. To address these issues, we construct a hierarchical hyperspectral crop dataset (H2Crop) by integrating 30m-resolution EnMAP hyperspectral data with Sentinel-2 time series. With over one million annotated field parcels organized in a four-tier crop taxonomy, H2Crop establishes a vital benchmark for fine-grained agricultural crop classification and hyperspectral image processing. We propose a dual-stream Transformer architecture that synergistically processes these modalities. It coordinates two specialized pathways: a spectral-spatial Transformer extracts fine-grained signatures from hyperspectral EnMAP data, while a temporal Swin Transformer extracts crop growth patterns from Sentinel-2 time series. The designed hierarchy classification heads with hierarchical fusion then simultaneously delivers multi-level classification across all taxonomic tiers. Experiments demonstrate that adding hyperspectral EnMAP data to Sentinel-2 time series yields a 4.2% average F1-scores improvement (peaking at 6.3%). Extensive comparisons also confirming our method’s higher accuracy over existing deep learning approaches for crop type classification and the consistent benefits of hyperspectral data across varying temporal windows and crop change scenarios. Codes and dataset will be available at https://github.com/flyakon/H2Crop and www.glass.hku.hk.
@@
Crop type classification, precision agriculture, remote sensing, deep learning, hyperspectral data, Sentinel-2 time series, fine-grained crops
@@ 1 Introduction
Accurate crop type classification serves as the cornerstone for global food security, agricultural policy formulation, and precision farming management. Satellite remote sensing has become a pivotal tool for large-scale agricultural monitoring due to its broad coverage and cost-effectiveness (Yuan et al., 2020; Weiss et al., 2020; Victor et al., 2024; Liang et al., 2024; Fang et al., 2024). The European Space Agency (ESA)’s Sentinel-2 satellites, with its 10-meter spatial resolution and multispectral bands, has significantly enhanced the ability to identify spatial details in crop classification (Blickensdörfer et al., 2022; Segarra et al., 2020). However, its spectral resolution limits its capability to differentiate crops with highly similar morphological and phenological traits.
@@
While machine learning (ML) methods have been widely adopted in crop mapping and achieved substantial progress, their limited model capacity and the inherent complexity of crop classification tasks restrict current applications to dominant staple crops (e.g., wheat, maize, rice). Most ML-based frameworks struggle to differentiate morphologically similar species or crops with overlapping phenological cycles, primarily due to insufficient nonlinear modeling capabilities.
@@
In scenarios where ground truth are derived from field surveys, crop classification samples typically correspond to time series of individual pixels rather than image sequences. This makes Long Short-Term Memory (LSTM) networks (Zhu et al., 2017; Ghojogh and Ghodsi, 2023) more suitable than Convolutional Neural Networks (CNNs) to model temporal dependencies for crop type classification. For instance, Rußwurm et al. (Rußwurm et al., 2023) designed an end-to-end LSTM framework for in-season crop identification, enabling early prediction of crop types during growing cycles. When processing spatiotemporal image sequences (e.g., multi-date satellite imagery), hybrid architectures combining CNNs and LSTMs have also gained traction: CNNs extract spatial-spectral features from individual images, while LSTMs model temporal evolution. Turkoglu et al. (Turkoglu et al., 2021) exemplified this paradigm, achieving fine-grained hierarchical crop classification through spatiotemporal feature fusion.
@@
H2Crop addresses three critical limitations in existing datasets: First, it surpasses hyperspectral-only dataset with large-scale annotations. Second, it extends conventional LPIS-derived crop datasets (e.g., ZueriCrop) through paired hyperspectral data. Third, it provides substantially greater spatial coverage and hierarchical granularity compared to other multi-modal datasets. The inclusion of parcel-based annotations and hyperspectral data enhances its utility for real-world agricultural monitoring applications, establishing H2Crop as an unprecedented resource for advancing fine-grained crop classification research.
@@ 2 Hierarchical Hyperspectral Crop (H2Crop) Dataset
The Hierarchical Hyperspectral Crop (H2Crop) represents the first dataset integrating time-series Sentinel-2 observations with EnMAP hyperspectral imagery for hierarchical fine-grained crop classification. It features a four-tier taxonomic system encompassing 6, 36, 82, and 101 crop types. Figure 1 illustrates the dataset’s spatial distribution across France and some representative samples, including Sentinel-2 data, exemplar crops from each hierarchical level (level 1 to level 4), and spectral curve derived from EnMAP data.
@@ 2.2 Dataset Construction Process
For dataset splitting, we implemented a frequency-aware partitioning method to handle the uneven spatial distribution of rare crops. After identifying the least frequent crop in each image as its signature crop, we sort all samples by signature crop frequency and allocated them to training (60%), validation (20%), and test (20%) sets to ensure all classes appear in each subset.
@@ 2.3 Overview of Dataset
Labels. The hierarchical labels provide four levels of crop classification granularity (6, 36, 82, and 101 crop types) at Sentinel-2’s 10m resolution. This multi-tier annotation system enables both coarse and fine-grained analysis.
@@
Figure 3 presents key statistical insights into the H2Crop dataset’s composition. Figure 3 (a) reveals the distribution of agricultural parcel sizes across the dataset’s 1,211,418 parcels, where most parcels (median area: 0.95 ha) fall below 2 hectares, though the mean area reaches 2.13 ha due to a long-tailed distribution of larger parcels. Figure 3 (b) quantifies the categorical complexity within individual samples, demonstrating that typical observation areas contain 7-15 distinct Level 4 crop types, reflecting exceptional biodiversity in agricultural landscapes. Figure 3 (c) employs logarithmic scaling to visualize the area distribution of the 19 most prevalent Level 4 crop types, with all remaining categories aggregated as ”Others.” This representation highlights the natural imbalance in crop cultivation patterns, where certain dominant species occupy disproportionately large areas compared to numerous specialty crops. Such imbalance presents significant challenges for rare crop identification, which our dataset’s comprehensive hierarchical labeling and spatial coverage are specifically designed to address.
@@ 3 Method
We present a dual-stream Transformer architecture designed to synergistically process hyperspectral and multispectral time-series data for precise crop classification, as illustrated in Figure 5. The hyperspectral branch employs a Spectral-Spatial Decoupled Vision Transformer that separately analyzes spectral signatures through spectral vision transformer and spatial patterns via spatial transformer. Simultaneously, the multispectral temporal transformer adapts Video Swin Transformer principles to jointly model temporal features and spatial features from Sentinel-2 monthly composites. These modality-specific features are aligned through pixel shuffle upsampling and integrated via convolutional layers, ensuring efficient cross-resolution fusion. The hierarchical classification heads comprises four cascaded prediction layers, each combining three information streams: (1) fused spectral-temporal features, (2) probability outputs from the preceding classification stage, and (3) historical crop type embeddings from the previous growing season. This design explicitly encodes agricultural knowledge while maintaining computational efficiency through single-convolution-layer heads at each hierarchy level.
@@ 3.2 Multispectral Temporal Transformer
3.3 Hierarchy Classification Heads
@@ 3.5 Validation
Since our main purpose of the proposed method in this paper is to verify the effectiveness of HSI data and time series MSI data, we organize subsequent experiments focus on the role of hyperspectral data in different occasion with only multispectral temporal data or adding prior crop type data. Our validation strategy systematically evaluates two core hypotheses: (1) hyperspectral data provides statistically significant improvements over multispectral-temporal baselines across diverse input configurations, and (2) the proposed hierarchical architecture outperforms conventional crop classification approaches. The metrics we adopt to verify include precision (P), recall (R) and F1 scores (F1):
@@ 4.2 Performance Comparison with Conventional Architectures
Further analysis reveals two key insights: First, our method dominates fine-grained classification (Levels 3-4), with Level 4 F1 reaching 49.6 (at least 3 points higher than alternatives), attributable to the Transformer’s capacity to capture subtle spectral variations. Second, while 3DCNN slightly outperforms other benchmarks at Levels 2-3, it still trails our method by 3-5 points at finer levels, indicating 3D convolutions’ limited efficacy for highly nuanced category differentiation.
@@ 4.3 Performance Comparison of Hierarchical Head and Independent Heads
The superiority of our hierarchical head manifests in two key aspects: First, joint training implicitly learns taxonomic dependencies (e.g., the logical chain ”cereals → wheat → winter wheat”), reducing classification conflicts. Second, confidence scores from higher-level predictions serve as priors for subordinate levels, enhancing discrimination of ambiguous samples. Notably, the slight advantage of independent heads at coarse levels (Levels 1-2) likely stems from their exclusive focus on single tasks, but this benefit diminishes with finer classification granularity (Levels 3-4 show F1 improvements of 0.2-1.9 percentage-points). Overall, our hierarchical head achieves significant gains in fine-grained classification and logical consistency at minimal cost to coarse-level performance (0.9 F1 reduction at Level 1), better aligning with real-world agricultural decision-making requiring multi-level reasoning.
@@ 5.1 Differential Impacts of Priors and Hyperspectral Data on Changed and Unchanged Crops
Table 6 details the hierarchical classification accuracy under four data configurations across changed and unchanged areas. The most striking finding reveals that in changed areas, introducing prior decreases average F1-score from 25.8 (S2-only) to 18.6 (S2+Prior), a 7.2 percentage-point degradation. Conversely, in unchanged areas, the same configuration elevates F1 from 31.1 to 68.0, a 36.9 percentage-point improvement. This extreme contrast demonstrates prior’s effectiveness is highly contingent upon crop planting continuity. Figure 11 further illustrates these results through matrix-style bar charts: Top row showing F1 improvements from hyperspectral data without (left) and with (right) prior knowledge; Bottom row showing F1 changes from prior knowledge without (left) and with (right) hyperspectral data. Key observations include: (1) hyper consistently delivers 5.1-6.4 percentage-point F1 gains in changed areas regardless of prior’s presence; (2) hyper’s improvements in unchanged areas are marginal (0.5-1.6 percentage-points); (3) hyper partially mitigates prior’s adverse effects in changed areas (S2+Prior+Hyper outperforms S2+Prior by 6.4 percentage-points).
@@
Mechanistically, when crop types change, prior misleads the network into reinforcing incorrect feature associations, causing severe accuracy deterioration. Hyperspectral data counteracts this by providing objective spectral evidence—evidenced by Level 4 F1 improvements of 7.0 percentage-points (S2+Prior+Hyper vs. S2+Prior) in changed areas. In unchanged regions, priors dominates classification performance while hyperspectral data serves primarily as a fine-tuning agent (all-level variations ¡1 percentage-point). Collectively, these results robustly demonstrate that hyper provides genuine discriminative information capable of substantially enhancing fine-grained crop classification, particularly in areas with frequent crop rotation where its contributions prove decisive. Our findings offer critical operational guidance: (1) in dynamic farming regions with frequent crop pattern changes, practitioners should prioritize hyperspectral data while avoiding prior; (2) in stable agricultural zones, prior can be selectively employed to boost performance, but must be coupled with hyperspectral data to mitigate potential negative impacts in transition areas.
@@ 5.3 Perspective and Future Directions
As a pioneering effort, we have validated the necessity of incorporating hyperspectral data for fine-grained crop classification and demonstrated the value of our dataset and methodology. Building upon these foundations, we identify three key directions for future advancement:
@@ 6 Conclusion
Comprehensive experiments systematically demonstrate the stable performance benefits from hyperspectral data across varying temporal spans, crop change scenarios, and model scales, with particularly notable improvements at fine-grained classification levels. These results conclusively establish the necessity of hyperspectral information for advanced crop differentiation.
The work’s significance manifests across multiple dimensions: For agricultural applications, the achieved high-precision classification capability directly supports precision farming, yield estimation, and food security monitoring. Methodologically, the revealed spectral-temporal complementarity principles provide new insights for refined agricultural analytics. The open-sourced H2Crop dataset will continue enabling advancements in crop identification, change detection, and anomaly monitoring research. Particularly relevant to climate change adaptation, our method effectively captures crop feature variations under emerging cultivation patterns, offering critical technical support for climate-smart agricultural monitoring systems. Future expansions of the dataset’s geographical coverage and crop diversity will further propel the transformation of agricultural remote sensing from extensive monitoring to intelligent, fine-grained observation systems.
Data Availability
Codes and dataset will be available on the website of the Jockey Club STEM Lab of Quantitative Remote Sensing, HKU (https://glass.hku.hk) and https://github.com/flyakon/H2Crop.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

# Fine-grained Hierarchical Crop Type Classification from Integrated Hyperspectral EnMAP Data and Multispectral Sentinel-2 Time Series: A Large-scale Dataset and Dual-stream Transformer Method

@@ Abstract
Fine-grained crop type classification serves as the fundamental basis for large-scale crop mapping and plays a vital role in ensuring food security. It requires simultaneous capture of both phenological dynamics (obtained from multi-temporal satellite data like Sentinel-2) and subtle spectral variations (demanding nanometer-scale spectral resolution from hyperspectral imagery). Research combining these two modalities remains scarce currently due to challenges in hyperspectral data acquisition and crop types annotation costs. To address these issues, we construct a hierarchical hyperspectral crop dataset (H2Crop) by integrating 30m-resolution EnMAP hyperspectral data with Sentinel-2 time series. With over one million annotated field parcels organized in a four-tier crop taxonomy, H2Crop establishes a vital benchmark for fine-grained agricultural crop classification and hyperspectral image processing. We propose a dual-stream Transformer architecture that synergistically processes these modalities. It coordinates two specialized pathways: a spectral-spatial Transformer extracts fine-grained signatures from hyperspectral EnMAP data, while a temporal Swin Transformer extracts crop growth patterns from Sentinel-2 time series. The designed hierarchical classification head with hierarchical fusion then simultaneously delivers multi-level crop type classification across all taxonomic tiers. Experiments demonstrate that adding hyperspectral EnMAP data to Sentinel-2 time series yields a 4.2% average F1-scores improvement (peaking at 6.3%). Extensive comparisons also confirm our method’s higher accuracy over existing deep learning approaches for crop type classification and the consistent benefits of hyperspectral data across varying temporal windows and crop change scenarios. Codes and dataset will be available at https://github.com/flyakon/H2Crop.
@@
Crop type classification, precision agriculture, remote sensing, deep learning, hyperspectral data, Sentinel-2 time series, fine-grained crop types
@@ 1 Introduction
Fine-grained crop type classification serves as the cornerstone for global food security, agricultural policy formulation, and precision farming management. Satellite remote sensing has become a pivotal tool for large-scale agricultural monitoring due to its broad coverage and cost-effectiveness (Yuan et al., 2020; Weiss et al., 2020; Victor et al., 2024; Liang et al., 2024; Fang et al., 2024). The European Space Agency (ESA)’s Sentinel-2 satellites, with their 10-meter spatial resolutions and multispectral bands, have significantly enhanced the ability to identify spatial details in fine-grained crop classification (Blickensdörfer et al., 2022; Segarra et al., 2020). However, its spectral resolution limits its capability to differentiate crops with highly similar morphological and phenological traits.
@@
While machine learning (ML) methods have been widely adopted in crop classification and mapping and achieved substantial progress, their limited model capacity and the inherent complexity of crop classification tasks restrict current applications to dominant staple crops (e.g., wheat, maize, rice). Most ML-based frameworks struggle to differentiate morphologically similar species or crops with overlapping phenological cycles, primarily due to insufficient nonlinear modeling capabilities.
@@
In scenarios where ground truth is derived from field surveys, crop classification samples typically correspond to time series of individual pixels rather than image sequences. This makes Long Short-Term Memory (LSTM) networks (Zhu et al., 2017; Ghojogh and Ghodsi, 2023) more suitable than Convolutional Neural Networks (CNNs) to model temporal dependencies for crop type classification. For instance, Rußwurm et al. (Rußwurm et al., 2023) designed an end-to-end LSTM framework for in-season crop identification, enabling early prediction of crop types during growing cycles. When processing spatiotemporal image sequences (e.g., multi-date satellite imagery), hybrid architectures combining CNNs and LSTMs have also gained traction: CNNs extract spatial-spectral features from individual images, while LSTMs model temporal evolution. Turkoglu et al. (Turkoglu et al., 2021) exemplified this paradigm, achieving fine-grained hierarchical crop classification through spatiotemporal feature fusion.
@@
H2Crop addresses three critical limitations in existing datasets. First, it surpasses hyperspectral-only dataset with large-scale annotations. Second, it extends conventional LPIS-derived crop datasets (e.g., ZueriCrop) through paired hyperspectral data. Third, it provides substantially greater spatial coverage and hierarchical granularity compared to other multi-modal datasets. The inclusion of parcel-based annotations and hyperspectral data enhances its utility for real-world agricultural monitoring applications, establishing H2Crop as an unprecedented resource for advancing fine-grained crop classification and mapping research.
@@ 2 Hierarchical Hyperspectral Crop (H2Crop) Dataset
The Hierarchical Hyperspectral Crop dataset (H2Crop) represents the first dataset integrating time-series Sentinel-2 observations with EnMAP hyperspectral imagery for hierarchical fine-grained crop classification. It features a four-tier taxonomic system encompassing 6, 36, 82, and 101 crop types. Figure 1 illustrates the dataset’s spatial distribution across France and some representative samples, including Sentinel-2 data, exemplar crops from each hierarchical level (level 1 to level 4), and spectral curve derived from EnMAP data.
@@ 2.2 Dataset Construction Process
For dataset splitting, we implement a frequency-aware partitioning method to handle the uneven spatial distribution of rare crops. After identifying the least frequent crop in each image as its signature crop, we sort all samples by signature crop frequency and allocate them to training (60%), validation (20%), and test (20%) sets to ensure all classes appear in each subset.
@@ 2.3 Overview of Dataset
Labels. The hierarchical labels provide four levels of crop classification granularity (6, 36, 82, and 101 crop types) at Sentinel-2’s 10m resolution. This multi-tier annotation system enables both coarse and fine-grained crop type classification and mapping research.
@@
Figure 3 presents key statistical insights into the H2Crop dataset’s composition. Figure 3 (a) reveals the distribution of agricultural parcel sizes across the dataset’s 1,211,418 parcels, where most parcels (median area: 0.95 ha) fall below 2 hectares, though the mean area reaches 2.13 ha due to a long-tailed distribution of larger parcels. Figure 3 (b) quantifies the categorical complexity within individual samples, demonstrating that typical observation areas contain 7-15 distinct Level 4 crop types, reflecting exceptional biodiversity in agricultural landscapes. Figure 3 (c) employs logarithmic scaling to visualize the area distribution of the 19 most prevalent Level 4 crop types, with all other categories aggregated as ”Others.” This representation highlights the natural imbalance in crop cultivation patterns, where certain dominant species occupy disproportionately large areas compared to numerous specialty crops. Such imbalance presents significant challenges for rare crop identification, which our dataset’s comprehensive hierarchical labeling and spatial coverage are specifically designed to address.
@@ 3 Method
We present a dual-stream Transformer architecture designed to synergistically process hyperspectral and multispectral time-series data for precise crop classification, as illustrated in Figure 5. The hyperspectral branch employs a Spectral-Spatial Decoupled Vision Transformer that separately analyzes spectral signatures through spectral vision transformer and spatial patterns via spatial transformer. Simultaneously, the multispectral temporal transformer adapts Video Swin Transformer principles to jointly model temporal features and spatial features from Sentinel-2 monthly composites. These modality-specific features are aligned through pixel shuffle upsampling and integrated via convolutional layers, ensuring efficient cross-resolution fusion. The hierarchical classification heads comprises four cascaded prediction layers, each combining three information streams: (1) fused spectral-temporal features, (2) probability outputs from the preceding classification head, and (3) historical crop type embeddings from the previous growing season. This design explicitly encodes agricultural knowledge while maintaining computational efficiency through single-convolution-layer heads at each hierarchy level.
@@ 3.2 Multispectral Temporal Transformer
3.3 Hierarchical Classification Heads
@@ 3.5 Validation
Since our main purpose in this paper is to verify the effectiveness of HSI data and time series MSI data, we organize subsequent experiments focus on the role of hyperspectral data in different occasion with only multispectral temporal data or adding prior crop type data. Our validation strategy systematically evaluates two core hypotheses: (1) hyperspectral data provides statistically significant improvements over multispectral-temporal baselines across diverse input configurations, and (2) the proposed hierarchical architecture outperforms conventional crop classification approaches. The metrics we adopt to verify include precision (P), recall (R) and F1 scores (F1):
@@ 4.2 Performance Comparison with Conventional Architectures
Further analysis reveals two key insights. First, our method dominates fine-grained classification (Levels 3-4), with Level 4 F1 reaching 49.6 (at least 3 points higher than alternatives), attributable to the Transformer’s capacity to capture subtle spectral variations. Second, while 3DCNN slightly outperforms other benchmarks at Levels 2-3, it still trails our method by 3-5 points at finer levels, indicating 3D convolutions’ limited efficacy for highly nuanced category differentiation.
@@ 4.3 Performance Comparison of Hierarchical Head and Independent Heads
The superiority of our hierarchical head manifests in two key aspects. First, joint training implicitly learns taxonomic dependencies (e.g., the logical chain ”cereals → wheat → winter wheat”), reducing classification conflicts. Second, confidence scores from higher-level predictions serve as priors for subordinate levels, enhancing discrimination of ambiguous samples. Notably, the slight advantage of independent heads at coarse levels (Levels 1-2) likely stems from their exclusive focus on single tasks, but this benefit diminishes with finer classification granularity (Levels 3-4 show F1 improvements of 0.2-1.9 percentage-points). Overall, our hierarchical head achieves significant gains in fine-grained classification and logical consistency at minimal cost to coarse-level performance (0.9 F1 reduction at Level 1), better aligning with real-world agricultural decision-making requiring multi-level reasoning.
@@ 5.1 Differential Impacts of Priors and Hyperspectral Data on Changed and Unchanged Crops
Table 6 details the hierarchical classification accuracy under four data configurations across changed and unchanged areas. The most striking finding reveals that in changed areas, introducing prior decreases average F1-score from 25.8 (S2-only) to 18.6 (S2+Prior), a 7.2 percentage-point degradation. Conversely, in unchanged areas, the same configuration elevates F1 from 31.1 to 68.0, a 36.9 percentage-point improvement. This extreme contrast demonstrates prior’s effectiveness is highly contingent upon crop planting continuity. Figure 11 further illustrates these results through matrix-style bar charts. Top row shows F1 improvements from hyperspectral data without (left) and with (right) prior knowledge. Bottom row shows F1 changes from prior knowledge without (left) and with (right) hyperspectral data. Key observations include: (1) hyper consistently delivers 5.1-6.4 percentage-point F1 gains in changed areas regardless of prior’s presence; (2) hyper’s improvements in unchanged areas are marginal (0.5-1.6 percentage-points); (3) hyper partially mitigates prior’s adverse effects in changed areas (S2+Prior+Hyper outperforms S2+Prior by 6.4 percentage-points).
@@
Mechanistically, when crop types change, prior misleads the network into reinforcing incorrect feature associations, causing severe accuracy deterioration. Hyperspectral data counteracts this by providing objective spectral evidence—evidenced by Level 4 F1 improvements of 7.0 percentage-points (S2+Prior+Hyper vs. S2+Prior) in changed areas. In unchanged regions, priors dominates classification performance while hyperspectral data serves primarily as a fine-tuning agent (all-level variations $\leq$1 percentage-point). Collectively, these results robustly demonstrate that hyper provides genuine discriminative information capable of substantially enhancing fine-grained crop classification, particularly in areas with frequent crop rotation where its contributions prove decisive. Our findings offer critical operational guidance: (1) in dynamic farming regions with frequent crop pattern changes, practitioners should prioritize hyperspectral data while avoiding prior; (2) in stable agricultural zones, prior can be selectively employed to boost performance, but must be coupled with hyperspectral data to mitigate potential negative impacts in transition areas.
@@ 5.3 Perspective and Future Directions
As a pioneering effort, we have validated the necessity of incorporating hyperspectral data for fine-grained crop classification and demonstrated the value of our dataset and methodology. Building upon these foundations, we identify three key directions for future advancement.
@@ 6 Conclusion
Comprehensive experiments systematically demonstrate the stable performance benefits from hyperspectral data across varying temporal spans, and crop change scenarios, with particularly notable improvements at fine-grained classification levels. These results conclusively establish the necessity of hyperspectral information for advanced crop differentiation.
The work’s significance manifests across multiple dimensions. For agricultural applications, the achieved high-precision classification capability directly supports precision farming, yield estimation, and food security monitoring. Methodologically, the revealed spectral-temporal complementarity principles provide new insights for refined agricultural analytics. The open-sourced H2Crop dataset will continue enabling advancements in crop identification and change detection. Particularly relevant to climate change adaptation, our method effectively captures crop feature variations under emerging cultivation patterns, offering critical technical support for climate-smart agricultural monitoring systems. Future expansions of the dataset’s geographical coverage and crop diversity will further propel the transformation of agricultural remote sensing from extensive monitoring to intelligent, fine-grained observation systems.